Find the following percentiles for the standard normal distribution. Interpolate where appropriate. (Round your answers to two decimal places.)

(a)    81st


(b)    19th


(c)    76th


(d)    24th


(e)    10th
Answer #1
Concepts and reason

A statistical measure used to indicate the value below certain percentage of observations in a specified group is called percentile.

For example, the 50th percentile is the middle value of the group which is under study. In other words the value represents the median value.

Fundamentals

The α\alpha % percentile is defined as follows:

P(Xx)=α%P\left( {X \le x} \right) = \alpha \%

The excel function to compute the percentile value is =NORMSINV(probability){\rm{ = NORMSINV(probability)}}

(a)

The mean and standard deviation of the standard normal distribution is μ=0\mu = 0 and standard deviation σ=1\sigma = 1

The 81st percentile is calculated as follows:

p81=NORMSINV(0.81)=0.8778960.88\begin{array}{c}\\{p_{81}} = NORMSINV\left( {0.81} \right)\\\\ = 0.877896\\\\ \approx 0.88\\\end{array}

(b)

The mean and standard deviation of the standard normal distribution is μ=0\mu = 0 and standard deviation σ=1\sigma = 1

The 19th percentile is calculated as follows:

p19=NORMSINV(0.19)=0.87790.88\begin{array}{c}\\{p_{19}} = NORMSINV\left( {0.19} \right)\\\\ = - 0.8779\\\\ \approx - 0.88\\\end{array}

(c)

The mean and standard deviation of the standard normal distribution is μ=0\mu = 0 and standard deviation σ=1\sigma = 1

The 76th percentile is calculated as follows:

p76=NORMSINV(0.76)=0.7063030.71\begin{array}{c}\\{p_{76}} = NORMSINV\left( {0.76} \right)\\\\ = 0.706303\\\\ \approx 0.71\\\end{array}

(d)

The mean and standard deviation of the standard normal distribution is μ=0\mu = 0 and standard deviation σ=1\sigma = 1

The 24th percentile is calculated as follows:

p24=NORMSINV(0.24)=0.7063030.71\begin{array}{c}\\{p_{24}} = NORMSINV\left( {0.24} \right)\\\\ = - 0.706303\\\\ \approx - 0.71\\\end{array}

(e)

The mean and standard deviation of the standard normal distribution is μ=0\mu = 0 and standard deviation σ=1\sigma = 1

The 10th percentile is calculated as follows:

p10=NORMSINV(0.10)=1.281551.28\begin{array}{c}\\{p_{10}} = NORMSINV\left( {0.10} \right)\\\\ = - 1.28155\\\\ \approx - 1.28\\\end{array}

Ans: Part a

The 81st percentile for the standard normal distribution is 0.88.

Part b

The 19th percentile for the standard normal distribution is -0.88.

Part c

The 76th percentile for the standard normal distribution is 0.71.

Part d

The 24th percentile for the standard normal distribution is -0.71.

Part e

The 10th percentile for the standard normal distribution is -1.28.